Operations Report – December 19th

Crew 185
Operations Report 12/19/2017
SOL: 1
Name of person filing report: A.Passeron
Non-nominal systems:
– Heater :
The heater was not working properly on Sunday, even after reseting it. Shannon adviced us to change the air filter. On Monday morning, Thibault and I removed the air filter. It had not been changed since April 2016. At 10am, Ilaria and David went to city to buy a battery for the generator and a pack of air filters. We replaced the filter at 3:00pm. The heater is now working well.
– Generator :
On Sunday, the generator had a failure at 8:00pm, Fault message reading “Low battery”. Jump-start from Shannon’s Shuttle battery did not solve the issue. Generator had a second failure at 9:00pm on sunday, so we definitely turned it off for the night.
On Monday, Shannon helped us to change the battery at 3:00pm. We then drained oil from the radiator by pouring clean cooling liquid. The oil of the generator was also very dirty. It was so sticky that we had to start the generator for a few minutes to heat the oil and thin it. We thus were able to remove the oil filter, but only a fraction of dirty oil came out of the generator. We tried to remove as much dirty oil as possible by filling the motor with clean oil, but a valve seemed to block the oil circulation. After pouring as much clean oil as we could, we replaced the oil filter. The generator was finally turned on at 5:00pm and runned properly for the whole night.
Today, Shannon called a mechanic to definitely clean the oil circuit and fix the generator.
– Spacesuit 2 : After the engineering EVA at 12:40pm, we noticed that the ventilation systems of spacesuits N°2 and N°3 were not working. At 3:00pm, Thibault and I opened the backpack of spacesuit N°2 to repair the ventilation system. The backpack of spacesuit N°2 was full of dust. We removed the dust and the fans are now operationnal.
– Spacesuit 3 : At 3:20pm, Thibault and I opened the backpack of spacesuit N°3 to repair the ventilation system. The left fan in the backpack of spacesuit N°3 was blocked. We carefully unjammed the fan and cleaned it with blow-off canned air. It is now operationnal.
Generator (hours run):
Generator turned off by Shannon at 8:50 am.
Generator turned on by Shannon at 3:10pm.
Solar— SOC @ 7:40pm : 100%
Diesel – 80%
Propane – 58 percent volume
Ethanol Free Gasoline (5 Gallon containers for ATV) – 3.5 Gallons
Water (trailer) – 150 gallons
Water (static) – 270 gallons
Trailer to Static Pump used – No
Static to Loft Pump used – Yes
Water Meter: 50 gallons
Toilet tank emptied: Yes
ATVs Used: No
Oil Added? Yes, all the ATVs have been filled with ethanol free gasoline and are full (Except the blue Yamaha 300 without a battery)
ATV Fuel Used:
# Hours the ATVs were Used today: 00:00 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Opportunity rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Curiosity rover used:  No
Hours:
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: Ventilation systems of spacesuits N°2 and N°3 repaired.
Summary of Hab operations: All nominal
Summary of GreenHab operations: All nominal
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support: